# Homomorphism, substructure and ideal: Elementary but rigorous aspects of renormalization group or hierarchical structure of topological orders

### v1

@@ VI Acknowledgement
YF thanks Guangyue Ji, Ha-Quang Trung, and Bo Yang for the interesting discussions on the hierarchical structure of fractional quantum Hall systems. YF also thanks Takamasa Ando, Weiguang Cao, and Pochung Chen for interesting discussions on noninvertible symmetries, Takumi Fukushima for sharing his knowledge of fracton topological orders, and Yifan Liu for reminding us the significance representing RG as a projection in a different research project and Chihiro Matsui and Hosho Katsura for sharing their knowledge on the partial solvable systems and for the related discussions. YF thanks Sylvain Ribault for many useful comments on the earlier version of the draft. YF thanks Gerard Watts for pointing out the correspondence between the nontrivial homomorphism from the tricritical Ising CFT to the Ising CFT and phase diagram in [237] and Jurgen Fuchs for notifying terminologies in category theory and for related useful comments. YF thanks support from NCTS and CTC.
@@ Appendix A Adding integral of motions to Hamiltonian or inducing gauge fixing
In this section, we propose an alternative method representing RG by implementing an analogy to gauge fixing in lattice gauge theory and the quotient by ideal. As in the main text, we assume the UV Hamiltonian $H_{(1)}$ has $\mathbf{A}_{(1)}$ symmetry and it can be deformed to Hamiltonian $H_{(2)}$ with quotient ring $\mathbf{A}_{(1)}/\text{Ker}\rho=\mathbf{A}_{(2)}$ symmetry. Then we take the linear basis of $\text{Ker}\rho$ as $\{s_{(1),i}\}_{i=1}^{n}$ and consider deformations of the Hamiltonian $H_{(1)}$ under some terms represented as $Q_{s_{(1),i}}$ where $Q$ is matrix representations of the ideal.
Under the RG flow $\rho$, $\text{Ker}\rho$ can be sent to zero and this is the most fundamental phenomenology in the main text. In other words, if this condition $\text{Ker}\rho=0$ can be enforced by adding some terms to the Hamiltonian $H_{(1)}$, one can expect the deformed Hamiltonian will capture the properties of the Hamiltonian $H_{(2)}$. In the main text, we interpret the condition $\text{Ker}\rho=0$ as mass or anyon condensation, but one can interpret this condition as another familiar phenomenon in physics. This can be interpreted as a gauge fixing.
@@
By taking respective high energy sectors labeled by $Q_{s_{(1),i}}=q_{(1),i}$ where $q_{(1),i}$ is an eigenvalue and there exists $i$ with $q_{(1),i}\neq 0$, one may expect to obtain a series of deformed ring structures $\mathbf{A}_{(1),\{s_{(1),i}=q_{i}\}}$ for each sectors. However, because of the property of $a_{(1)}\times\text{Ker}\rho=\text{Ker}\rho$ for all $a_{(1)}\in\mathbf{A}_{(1)}$, the structure $\{q_{(1),i}\}$ completely determines all $a_{(1)}\in\mathbf{A}_{(1)}$ as a scalar. In other words, only the sector $\text{Ker}\rho=0$ generates a nontrivial emergent ring structure or symmetry denoted as $\mathbf{A}_{(1)}/\text{Ker}\rho$.
We also note the analogy between Eq. (59) and effective Hamiltonian appearing in generalized Gibbs ensamble[260, 261] which has a close connection to the quantum Korteweg–De Vries equation[262] (For more general aspects, see discussions in [209]). In other words, the perturbations added to $H_{(1)}$ induce thermalization, and one can interprete RG and resultant $c$- or $g$-theorem as a law of thermodynamics[253, 254, 255]. To some extent, this is a standard phenomenological understanding of RG, but one can see the analogy more evidently by observing the structures coming from the ideal.
@@ Appendix B Conjecture on coset representation of symmetry preserving domain walls
Generalizing the argument in existing literature, we conjecture the decomposition of coset CFTs[107, 108, 109, 110]. The symbol “///" in this section is different from that of the quotient ring in the main text, so we placed the arguments here. As far as we have checked, the construction of ideals in the main text has not been achieved in general, and the study of the possible homomorphism between theories described by coset CFTs will be useful. The correspondence between coset conformal field theory and massless RG flow has been studied in [111] and its connection to current-current deformation has been studied in the context of $\lambda$ deformation[263, 264, 265, 266]. Related quantum wire junction problems have been studied in [267, 268, 269]. We also note more recent literature from different research motivations[270, 271]. In this section, we denote chiral or antichiral CFT as $M$ or $\overline{M}$ and label them by lower index IR or UV to identify them as UV and IR theories respectively.
In $Z_{N}$ symmetric models, there exist two types of $Z_{N}$ symmetry preserving RG flow, anomaly matching flow and anomaly cancellation flow[67, 69]. For example, in the flow between $SU(N)_{k}$ WZW model to $SU(N)_{k^{\prime}}$ WZW model, the former corresponds to the periodicity $k-k^{\prime}\ (\text{mod}.N)$ and the latter produces the periodicity $k+k^{\prime}\ (\text{mod}.N)$. The theory $M_{\text{UV}}\otimes\overline{M_{\text{IR}}}$ or $M_{\text{UV}}\otimes M_{\text{IR}}$ has nonchiral or chiral (half) integer spin simple current corresponding to anomaly matching or anomaly cancellations. For anomaly matching flow, we propose the following coset decomposition or Witt equivalence,
@@
For anomaly cancellation flow, we also propose the following Witt equivalent relation,
@@
Figure 5: Domain wall or tri-wire junction problem corresponding to the massless RG. The fields $\varphi$ and $\overline{\varphi}$ are chiral or antichiral fields respectively. An exact description of the theory appearing in the domain wall is an open problem in general, but there exist respective researches verifying this picture (anomaly matching case, in particular)[267, 268, 269].
For the $Z_{N}$ symmetric model, the (half) integer spin simple current should govern the classification of $Z_{N}$ symmetry preserving domain walls. As a straightforward generalization, nonsimple current[274] with integer spin, which can be called integer spin nonsimple current, should govern the anomaly classification of noninvertible symmetric models and this is related to maverick coset in the literature[275, 276, 277, 278]. Because of its noninvertible nature, it seems necessary to study the Deligne product structure or chiral-chiral chiral-antichiral anyon condensations of such theories more carefully. Related to this direction, it is also known that, if two systems have the same fusion rule, particular types of modular invariants appear[279, 280].

### v2

@@ VI Acknowledgement
YF thanks Guangyue Ji, Ha-Quang Trung, and Bo Yang for the interesting discussions on the hierarchical structure of fractional quantum Hall systems. YF also thanks Takamasa Ando, Weiguang Cao, and Pochung Chen for interesting discussions on noninvertible symmetries, Takumi Fukushima for sharing his knowledge of fracton topological orders, and Yifan Liu for reminding us of the significance of representing RG as a projection in a different research project, and Chihiro Matsui and Hosho Katsura for sharing their knowledge on the partial solvable systems and for the related discussions. YF thanks Sylvain Ribault for many useful comments on the earlier version of the draft. YF thanks Gerard Watts for pointing out the correspondence between the nontrivial homomorphism from the tricritical Ising CFT to the Ising CFT and phase diagram in [240] and Jurgen Fuchs for notifying us of terminology in category theory and for related useful comments. YF thanks the support from NCTS and CTC.
@@ Appendix A Adding integral of motions to Hamiltonian or inducing gauge fixing
In this section, we propose an alternative method for representing RG by implementing an analogy to gauge fixing in lattice gauge theory and the quotient by an ideal. As in the main text, we assume the UV Hamiltonian $H_{(1)}$ has $\mathbf{A}_{(1)}$ symmetry and it can be deformed to Hamiltonian $H_{(2)}$ with quotient ring $\mathbf{A}_{(1)}/\text{Ker}\rho=\mathbf{A}_{(2)}$ symmetry. Then we take the linear basis of $\text{Ker}\rho$ as $\{s_{(1),i}\}_{i=1}^{n}$ and consider deformations of the Hamiltonian $H_{(1)}$ under some terms represented as $Q_{s_{(1),i}}$ where $Q$ is matrix representations of the ideal.
Under the RG flow $\rho$, $\text{Ker}\rho$ can be sent to zero, and this is the most fundamental phenomenology in the main text. In other words, if this condition $\text{Ker}\rho=0$ can be enforced by adding some terms to the Hamiltonian $H_{(1)}$, one can expect the deformed Hamiltonian to capture the properties of the Hamiltonian $H_{(2)}$. In the main text, we interpret the condition $\text{Ker}\rho=0$ as mass or anyon condensation, but one can interpret this condition as another familiar phenomenon in physics. This can be interpreted as a gauge fixing.
@@
By taking respective high energy sectors labeled by $Q_{s_{(1),i}}=q_{(1),i}$ where $q_{(1),i}$ is an eigenvalue and there exists $i$ with $q_{(1),i}\neq 0$, one may expect to obtain a series of deformed ring structures $\mathbf{A}_{(1),\{s_{(1),i}=q_{i}\}}$ for each sectors. However, because of the property of $a_{(1)}\times\text{Ker}\rho=\text{Ker}\rho$ for all $a_{(1)}\in\mathbf{A}_{(1)}$, the structure $\{q_{(1),i}\}$ completely determines all $a_{(1)}\in\mathbf{A}_{(1)}$ as a scalar. In other words, only the sector $\text{Ker}\rho=0$ generates a nontrivial emergent ring structure or symmetry denoted as $\mathbf{A}_{(1)}/\text{Ker}\rho$ when fixing or condensing the structure $\text{Ker}\rho$. However, by taking a subrng of $\mathbf{R}=\{r_{i}\}_{i=1}^{m}\subset\text{Ker}\rho$ with $m<n$, one can still obtain a nontrivial deformed ring when the subrng $\mathbf{R}$ does not form an ideal in $\mathbf{A}_{1}$. Hence, we propose the following structure as a deformed ring (or ring after projection),
$\mathbf{A}_{(1),\{r_{(1),i}=q_{i}\}},\ \text{with}\ \mathbf{R}=\{r_{i}\}_{i=1}^{m}\subset\text{Ker}\rho$ (64)
The value $q_{i}$ corresponds to the quantum dimension of the objects as a consequence of homomorphism between $\mathbf{A}_{(1)}\rightarrow\mathbf{C}$ where $\mathbf{C}$ is the complex number fields, but this will require more detailed investigations.
We also note the analogy between Eq. (62) and the effective Hamiltonian appearing in generalized Gibbs ensemble[265, 266] which has a close connection to the quantum Korteweg–De Vries equation[267] (For more general aspects, see discussions in [212]). In other words, the perturbations added to $H_{(1)}$ induce thermalization, and one can interprete RG and resultant $c$- or $g$-theorem as a law of thermodynamics[258, 259, 260]. To some extent, this is a standard phenomenological understanding of RG, but one can see the analogy more evidently by observing the structures coming from the ideal.
@@ Appendix B Conjecture on coset representation of symmetry preserving domain walls
Generalizing the argument in existing literature, we conjecture the decomposition of coset CFTs[106, 107, 108, 109]. The symbol “//" in this section is different from that of the quotient ring in the main text, so we placed the arguments here. As far as we have checked, the construction of ideals in the main text has not been achieved in general, and the study of the possible homomorphism between theories described by coset CFTs will be useful. The correspondence between coset conformal field theory and massless RG flow has been studied in [110], and its connection to current-current deformation has been studied in the context of $\lambda$ deformation[268, 269, 270, 271]. Related quantum wire junction problems have been studied in [272, 273, 274]. We also note more recent literature from different research motivations[275, 276]. In this section, we denote chiral or antichiral CFT as $M$ or $\overline{M}$ and label them by lower index IR or UV to identify them as UV and IR theories, respectively.
In $Z_{N}$ symmetric models, there exist two types of $Z_{N}$ symmetry preserving RG flow, anomaly matching flow, and anomaly cancellation flow[66, 68]. For example, in the flow between $SU(N)_{k}$ WZW model to $SU(N)_{k^{\prime}}$ WZW model, the former corresponds to the periodicity $k-k^{\prime}\ (\text{mod}.N)$ and the latter produces the periodicity $k+k^{\prime}\ (\text{mod}.N)$. The theory $M_{\text{UV}}\otimes\overline{M_{\text{IR}}}$ or $M_{\text{UV}}\otimes M_{\text{IR}}$ has nonchiral or chiral (half) integer spin simple current corresponding to anomaly matching or anomaly cancellations. For anomaly matching flow, we propose the following coset decomposition or Witt equivalence,
@@
For the anomaly cancellation flow, we also propose the following Witt equivalent relation,
@@
Figure 5: Domain wall or tri-wire junction problem corresponding to the massless RG. The fields $\varphi$ and $\overline{\varphi}$ are chiral or antichiral fields, respectively. An exact description of the theory appearing in the domain wall is an open problem in general, but there exists respective research verifying this picture (anomaly matching case, in particular)[272, 273, 274].
For the $Z_{N}$ symmetric model, the (half) integer spin simple current should govern the classification of $Z_{N}$ symmetry-preserving domain walls. As a straightforward generalization, nonsimple current[279] with integer spin, which can be called integer spin nonsimple current, should govern the anomaly classification of noninvertible symmetric models, and this is related to maverick coset in the literature[280, 281, 282, 283]. Because of its noninvertible nature, it seems necessary to study the Deligne product structure or chiral-chiral chiral-antichiral anyon condensations of such theories more carefully. Related to this direction, it is also known that, if two systems have the same fusion rule, particular types of modular invariants appear[284, 285].